You serve a volleyball that has a mass of 2.1 kg. With what speed does the ball leave your hand if it has 945 J of kinetic energy? 

 



1
Expert's answer
2021-06-07T03:45:07-0400

"KE=mv^2/2\to v=\sqrt{2\cdot KE/m}=\sqrt{2\cdot945/2.1}=30\ (m/s)" . Answer